# The DECADE cosmic shear project III: validation of analysis pipeline using spatially inhomogeneous data

### v1

@@ Abstract
We present the pipeline for the cosmic shear analysis of the Dark Energy Camera All Data Everywhere (DECADE) weak lensing dataset: a catalog consisting of 107 million galaxies observed by the Dark Energy Camera (DECam) in the northern Galactic cap. The catalog derives from a large number of disparate observing programs and is therefore more inhomogeneous across the sky compared to existing lensing surveys. First, we use simulated data-vectors to show the sensitivity of our constraints to different analysis choices in our inference pipeline, including sensitivity to residual systematics. Next we use simulations to validate our covariance modeling for inhomogeneous datasets. Finally, we show that our choices in the end-to-end cosmic shear pipeline are robust against inhomogeneities in the survey, by extracting relative shifts in the cosmology constraints across different subsets of the footprint/catalog and showing they are all consistent within $1\sigma$ to $2\sigma$. This is done for forty-six subsets of the data and is carried out in a fully consistent manner: for each subset of the data, we re-derive the photometric redshift estimates, shear calibrations, survey transfer functions, the data vector, measurement covariance, and finally, the cosmological constraints. Our results show that existing analysis methods for weak lensing cosmology can be fairly resilient towards inhomogeneous datasets. This also motivates exploring a wider range of image data for pursuing such cosmological constraints.
@@ 1 Introduction
Over the past two decades, weak gravitational lensing (also referred to as weak lensing or cosmic shear) has emerged as a leading probe in constraining the cosmological parameters of our Universe (Asgari & Lin et al., 2021; Secco & Samuroff & Samuroff et al., 2022; Amon & Gruen et al., 2022; Dalal & Li et al., 2023). Weak lensing refers to the subtle bending of light from distant “source galaxies” due to the large-scale matter distribution between the source and the observer (Bartelmann & Schneider 2001). Thus, weak lensing, through its sensitivity to the matter distribution, probes the large-scale structure (LSS) of our Universe and any processes that impact this structure; including cosmological processes such as modified gravity (e.g., Schmidt 2008) and primordial signatures (e.g., Anbajagane et al. 2024c; Goldstein et al. 2024), as well as a wide variety of astrophysical processes (e.g., Chisari et al. 2018; Schneider et al. 2019; Aricò et al. 2021; Grandis et al. 2024; Bigwood et al. 2024). Weak lensing has many novel advantages in the landscape of cosmological probes, the primary of which is that it is an unbiased tracer of the density field — unlike other tracers, such as galaxies — and does not require modeling or marginalizing over an associated bias parameter (Bartelmann & Schneider 2001). For these reasons, it is one of the leading probes of cosmology and has delivered some of our best constraints on cosmological parameters.
@@
This work serves three, key purposes. First, to detail the modeling/methodology choices of the cosmic shear analysis, and the robustness of our results to said choices. Second, to build on the null-tests of Paper I and show that our data vector (and cosmology) are not susceptible to contamination from systematic effects, such as correlated errors in the point-spread function (PSF) modeling. Finally, we test the impact of spatial inhomogeneity in the entire end-to-end pipeline used to extract the cosmology constraints. As highlighted in both Paper I and Paper II, the DECADE dataset contains some unique characteristics relative to other WL datasets; particularly, the spatial inhomogeneity in the image data coming from this dataset’s origin as an amalgamation of many different public observing programs. We perform a suite of tests where we rerun the end-to-end pipeline for different subsets of our data — where each subset contains specific kinds of galaxies (red/blue, faint/bright etc.) or contains objects measured in regions of the sky with better/worse image quality (changes in seeing, airmass, interstellar extinction etc.) — and show that our cosmology constraints are robust across such subsets.
This paper is structured as follows. In Section 2, we briefly describe the DECADE shape catalog, and in Section 3, we present the cosmology model used in the DECADE cosmic shear project. In Section 4, we outline the different components required for parameter inference, including our analytic covariance matrix. In Section 5, we check the robustness of our constraints across modeling choice in simulated data vectors. Section 6 details our tests on the sensitivity of our parameter constraints to spatial inhomoegenity and to different selections of the source galaxy catalog. We conclude in Section 7.
@@ 2 The DECADE dataset
The DECADE dataset is a galaxy shape catalog of 107 million galaxies, spanning $5,\!412\deg^{2}$ of the sky. The catalog is introduced in Paper I, alongside a suite of null-tests and shear calibrations made using image simulations of the survey data. The galaxy shears are estimated using the Metacalibration method (Sheldon & Huff 2017), with an approach designed to mimic that of DES Y3. In Paper II, we split the catalog into four tomographic bins and estimate the redshift distribution of the ensembles in each bin using the self-organizing maps photo-z (SOMPZ) method (Buchs et al. 2019; Myles & Alarcon et al. 2021), including cross-checking these estimates with alternative methods such as clustering redshifts (Schneider et al. 2006; Newman 2008; Ménard et al. 2013; Davis et al. 2017; Gatti & Vielzeuf et al. 2018; Gatti & Giannini et al. 2022).
@@
The DECADE dataset is constructed from available community data (collected up to December 2022) from the Dark Energy Camera (DECam, Flaugher et al. 2015) within the footprint of interest to us; see Figure 2 in Paper I for the survey footprint, and also Section 2.1 in that work. The nature of this dataset — as an amalgamation of available archival data, rather than as a dedicated cosmological survey program — results in significant inhomogeneities in the survey observing conditions (such as depth). This, in turn, propagates into the observed object properties and therefore, the detection/selection functions of galaxy samples. While such inhomogeneities in observing conditions exist for other photometric galaxy surveys, such as DES, the variations are smaller relative to the DECADE inhomogeneities; see Figure 1. The latter part of this work (Section 6) explicitly shows the consistency in cosmology constraints when limiting our dataset to different regions of the sky with more/less spatial inhomogeneity in the chosen image quality.
@@
In general, the constraining power of our catalog is fairly similar to that of DES Y3 (Paper I). We have a source-galaxy number density that is $\approx\!20\%$ lower than that of DES Y3 ($n_{g}=4.6\,\,\text{arcmin}^{-2}$ compared to $n_{g}=5.6\,\,\text{arcmin}^{-2}$); however, our footprint is $\approx\!25\%$ larger ($5412\deg^{2}$ vs. $4143\deg^{2}$). Thus, our lower number density is compensated by our larger area, and our shape catalog’s statistical precision is similar to that of DES Y3. Note, however, that our sample has slightly lower redshift ($\Delta\langle z\rangle\approx 0.02$), particularly in the highest tomographic bins (Paper II).
[…]
@@
We use the two-point correlation function of galaxy shapes, $\xi_{\pm}(\theta)$ (Bartelmann & Schneider 2001), as our primary cosmological measurement. We have
@@
×$\times\left[C_{EE}^{ij}(\ell)\pm C_{BB}^{ij}(\ell)\right],$ (1)
where the functions $G^{\pm}_{\ell}(x)$ are computed from Legendre polynomials $P_{\ell}(x)$ and averaged over angular bins (Krause et al. 2021). The $i$ and $j$ indices specify the two tomographic redshift bins from which the correlation function is calculated. Assuming the extended Limber approximation (Limber 1953; Loverde & Afshordi 2008) and in a spatially flat universe, the lensing power spectrum can be written as
[…]
where $\chi$ is the radial comoving distance, $\chi_{H}$ is the distance to the horizon, $P_{\text{NL}}$ is the nonlinear matter power spectrum, and $q(\chi)$ is the lensing efficiency defined via
[…]
where $\Omega_{\rm m}$ is the matter density today, $H_{0}$ is the Hubble parameter today, $a$ is the scale factor, and $n_{i}(\chi)$ is the redshift distribution of the bin $i$. The term $C^{ij}_{BB}$ will be non-zero due to the presence of intrinsic alignments (IA); [^1] In practice, lensing also generates a $C^{ij}_{BB}$ term but its amplitude is orders of magnitude lower than the $C^{ij}_{EE}$ term and is therefore ignored (e.g., Krause & Hirata 2010). see Section 3.2 for discussions on our IA model.
@@ 3.1 Nonlinear power spectrum
On small scales, the nonlinearity of structure formation as induced by gravity necessitates a modification of the matter power spectrum from the linear model, where the latter is typically calculated via a Boltzmann solver such as CAMB (Lewis et al. 2000) or Class (Lesgourgues 2011). Over the years, as the resolution and fidelity of N-body simulations have improved and so has the prescription for the nonlinear power spectrum. In the DES Y1 and Y3 cosmic shear analysis, HaloFit (Takahashi et al. 2012) was used and validated as the default nonlinear power spectrum. However, the study and joint-analysis between DES and the Kilo Degree Survey (KiDS) in DES and KiDS Collaborations et al. (2023) show that the updated version of HMCode (HMCode2020, Mead et al. 2020, 2021) provides a more accurate description when tested against the Euclid emulator (Euclid Collaboration: Knabenhans et al. 2019). At $k<10h{\rm Mpc}^{-1}$, HMCode2020 has an accuracy of $<2.5\%$. As a result, we adopt HMCode2020 as the fiducial model for our nonlinear power spectrum. Note that for the constraining power of the DECADE data, we do not expect any significant differences in constraints when switching between the HaloFit or HMCode2020 models. We explicitly test this modeling choice in our simulated chains below, and also check it directly on the data in Paper IV.
@@
The structure on these small scales is also impacted by the non-gravitational processes of baryons; most prominently, the ejection of gas from the halo due to feedback processes (Chisari et al. 2018). Hydrodynamical simulations (see Vogelsberger et al. 2020, for a review), which model these effects through approximate “subgrid” prescriptions, generate a variety of predictions for the properties of halos in simulations (e.g., Anbajagane et al. 2020; Lim et al. 2021; Lee et al. 2022; Stiskalek et al. 2022; Anbajagane et al. 2022a, b; Shao et al. 2023; Shao & Anbajagane 2024; Gebhardt et al. 2023) and therefore also for the changes in the nonlinear matter power spectrum (e.g., Chisari et al. 2018; Amon & Efstathiou 2022). These variations have been captured through phenomenological, halo-based models such as “Baryonification” (Schneider & Teyssier 2015; Schneider et al. 2019; Aricò et al. 2021; Anbajagane et al. 2024a) and the HMCode2020 model discussed above.
@@
The HMCode2020 model includes a $T_{\rm AGN}$ parameter that quantifies the baryonic contribution to the nonlinear matter power spectrum, and this parameter was varied in the joint-analysis of DES and KiDS (DES and KiDS Collaborations et al. 2023). In this work, our fiducial analysis does not use any of these techniques and follows DES Y3 (Secco & Samuroff et al. 2022; Amon et al. 2022): we remove any scales where the baryonic effects are significant and analyze the remaining scales with a model that does not account for baryons. The identification of scales to be removed is done using the procedure described in Section 4.1. However, we also perform a variant analysis in Paper IV that uses all measurements (i.e., does not perform any scale cuts) and instead varies the $T_{\rm AGN}$ parameter to include the effects of baryons.
@@ 3.2 Intrinsic alignment (IA)
While cosmic shear manifests in the spatial correlations of galaxy shapes, these shapes can also be spatially correlated in the absence of lensing due to the gravitational environment, or tidal fields, that the galaxies form within (see, e.g., the review Troxel & Ishak 2015, and references therein). This effect is referred to as intrinsic alignment.
The cosmic shear analysis of DES Y1 uses the nonlinear alignment model (NLA, Hirata & Seljak 2004; Bridle & King 2007; Joachimi et al. 2011). This model assumes that the IA power spectrum is proportional to the nonlinear matter power spectrum in a scale-independent way, through a function, $F[\chi(z)]$, given by
$F[\chi(z)]=A_{\rm IA}C_{1}\rho_{\rm crit}\frac{\Omega_{\rm m}}{D_{+}(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta},$ (4)
where $A_{\rm IA}$ is the amplitude parameter, $C_{1}=5\times 10^{-14}h^{-2}M_{\odot}^{-1}\text{Mpc}^{3}$, $\rho_{\rm crit}$ is the critical density at $z=0$, and $D_{+}(z)$ is the linear growth factor normalized to unity at $z=0$ (Bridle & King 2007).
The DES Y3 analysis uses the more general, Tidal Alignment and Tidal Torquing model (TATT, Blazek et al. 2019). The TATT model extends from the NLA model and includes a tidal torquing alignment mechanism. It allows for five free nuisance parameters: a tidal alignment amplitude ($a_{1}$) with redshift evolution ($\eta_{1}$); a tidal torquing amplitude ($a_{2}$) with redshift evolution ($\eta_{2}$); and a linear bias amplitude ($b_{\rm TA}$). The TATT model reduces to the NLA model when setting $a_{1}=A_{\rm IA}$, $\eta=\eta$, and $a_{2}=\eta_{2}=b_{\rm TA}=0$. Different analyses of DES Y3 have used either TATT (Secco & Samuroff et al. 2022; Amon et al. 2022; Samuroff et al. 2023) or NLA (DES and KiDS Collaborations et al. 2023) as their fiducial choice. For the DECADE cosmic shear analysis, given the unique nature of our dataset, we will adopt the more conservative TATT model as our fiducial choice but test the change in our results if we revert to the simpler NLA models. [^2] Given TATT has more free parameters, there is now a higher chance for the constraints to explore unexpected parts of parameter space, due to the model having more freedom than can be constrained with the data. For this reason, it is often useful to check our results using NLA as well.

### v2

@@ Abstract
We present the pipeline for the cosmic shear analysis of the Dark Energy Camera All Data Everywhere (DECADE) weak lensing dataset: a catalog consisting of 107 million galaxies observed by the Dark Energy Camera (DECam) in the northern Galactic cap. The catalog derives from a large number of disparate observing programs and is therefore more inhomogeneous across the sky compared to existing lensing surveys. First, we use simulated data-vectors to show the sensitivity of our constraints to different analysis choices in our inference pipeline, including sensitivity to residual systematics. Next we use simulations to validate our covariance modeling for inhomogeneous datasets. Finally, we show that our choices in the end-to-end cosmic shear pipeline are robust against inhomogeneities in the survey, by extracting relative shifts in the cosmology constraints across different subsets of the footprint/catalog and showing they are all consistent within $1\sigma$ to $2\sigma$. This is done for forty-six subsets of the data and is carried out in a fully consistent manner: for each subset of the data, we rederive the photometric redshift estimates, shear calibrations, survey transfer functions, the data vector, measurement covariance, and finally, the cosmological constraints. Our results show that existing analysis methods for weak lensing cosmology can be fairly resilient towards inhomogeneous datasets. This also motivates exploring a wider range of image data for pursuing such cosmological constraints.
@@ 1 Introduction
Over the past two decades, weak gravitational lensing (also referred to as weak lensing or cosmic shear) has emerged as a leading probe in constraining the cosmological parameters of our Universe (Asgari et al. 2021; Secco & Samuroff et al. 2022; Amon et al. 2022; Dalal et al. 2023; Wright et al. 2025). Weak lensing refers to the subtle bending of light from distant “source galaxies” due to the large-scale matter distribution between the source and the observer (Bartelmann & Schneider 2001). Thus, weak lensing, through its sensitivity to the matter distribution, probes the large-scale structure (LSS) of our Universe and any phenomenon that impact this structure; including cosmological ones such as modified gravity (e.g., Schmidt 2008) and primordial signatures (e.g., Anbajagane et al. 2024c; Goldstein et al. 2024; Anbajagane & Lee 2025a, b), as well as a wide variety of astrophysical ones (e.g., Chisari et al. 2018; Schneider et al. 2019; Aricò et al. 2021; Grandis et al. 2024; Bigwood et al. 2024). Weak lensing has many advantages in the landscape of cosmological probes, the primary of which is that it is an unbiased tracer of the density field — unlike other tracers, such as galaxies — and does not require modeling or marginalizing over an associated bias parameter (Bartelmann & Schneider 2001). Since the lensing signal is also sensitive to the distance between the source galaxy, the intervening matter, and the observer, lensing is also a probe of the geometry of our Universe. For these reasons, it is one of the leading probes of cosmology and has delivered some of our best constraints on cosmological parameters.
@@
This work serves three, key purposes. First, to detail the modeling/methodology choices of the cosmic shear analysis, and the robustness of our results to said choices. Second, to build on the null-tests of Paper I and show that our data vector (and cosmology) are not susceptible to contamination from systematic effects, such as correlated errors in the point-spread function (PSF) modeling. Finally, we show that the end-to-end cosmology inference pipeline is resilient to the impact of spatial inhomogeneities. As highlighted in both Paper I and Paper II, the DECADE dataset contains some unique characteristics relative to other weak lensing datasets; particularly, the spatial inhomogeneity in the image data coming from this dataset’s origin as an amalgamation of many different public observing programs. We perform a suite of tests where we rerun the end-to-end pipeline for different subsets of our data — where each subset contains specific kinds of galaxies (red/blue, faint/bright, etc.) or contains objects measured in regions of the sky with better/worse image quality (changes in seeing, airmass, interstellar extinction, etc.) — and show that our cosmology constraints are robust across such subsets.
This paper is structured as follows. In Section 2, we briefly describe the DECADE shape catalog, and in Section 3, we present the cosmology model used in the DECADE cosmic shear project. In Section 4, we outline the different components required for parameter inference, including our analytic covariance matrix. In Section 5, we check the robustness of our constraints across modeling choice in simulated data vectors. Section 6 details our tests on the sensitivity of our parameter constraints to spatial inhomogeneity and to different selections of the source galaxy catalog. We conclude in Section 7.
@@ 2 The DECADE dataset
The DECADE dataset is a galaxy shape catalog of 107 million galaxies, spanning $5,\!412\deg^{2}$ of the sky. The catalog is introduced in Paper I, alongside a suite of null-tests and shear calibrations made using image simulations of the survey data. The galaxy shears are estimated using the Metacalibration method (Sheldon & Huff 2017), with an approach designed to mimic that of DES Y3 (Gatti & Sheldon et al. 2021). In Paper II, we split the catalog into four tomographic bins and estimate the redshift distribution of the ensembles in each bin using the self-organizing maps photometric redshifts (SOMPZ) method (Buchs et al. 2019; Myles & Alarcon et al. 2021), including cross-checking these estimates with alternative methods such as clustering redshifts (Schneider et al. 2006; Newman 2008; Ménard et al. 2013; Davis et al. 2017; Gatti & Vielzeuf et al. 2018; Gatti & Giannini et al. 2022).
@@
The DECADE dataset is constructed from available community data (collected up to December 2022) from the Dark Energy Camera (DECam, Flaugher et al. 2015) within the footprint of interest to us; see Figure 2 in Paper I for the survey footprint, and also Section 2.1 in that work. The nature of this dataset — as an amalgamation of available archival data, rather than as a dedicated cosmological survey program — results in significant inhomogeneities in the survey observing conditions (such as depth). This, in turn, propagates into the observed object properties and therefore, the detection/selection functions of galaxy samples. While such inhomogeneities in observing conditions exist for other photometric galaxy surveys, such as DES, the variations are smaller relative to the DECADE inhomogeneities; see Figure 1 (which is discussed in more detail in Section 4.2). The latter part of this work (Section 6) explicitly shows the consistency in cosmology constraints when limiting our dataset to different regions of the sky with more/less spatial inhomogeneity in the chosen image quality.
@@
In general, the constraining power of our catalog is fairly similar to that of DES Y3 (Paper I). We have a source-galaxy number density that is $\approx\!20\%$ lower than that of DES Y3 ($n_{\rm eff}=4.6\,\,\text{arcmin}^{-2}$ for DECADE compared to $n_{\rm eff}=5.6\,\,\text{arcmin}^{-2}$ in DES Y3111Here, $n_{\rm eff}$ is the effective source-galaxy number density that takes into account the galaxy weight. See Paper I for more details on how these weights are generated.); however, our footprint is $\approx\!25\%$ larger ($5412\deg^{2}$ vs. $4143\deg^{2}$). Thus, our lower number density is compensated by our larger area, and our shape catalog’s statistical precision is similar to that of DES Y3 (Gatti & Sheldon et al. 2021). Note, however, that our sample has a slightly lower mean redshift ($\Delta\langle z\rangle\approx 0.02$), particularly in the highest tomographic bins (Paper II).
Figure 1: A set of maps from the DECADE (top) and DES Y3 (bottom) datasets. From left to right, we show the (apparent) magnitude limit map in the $i$-band (estimated at $S/N=10$ within a $\arcsec2\arcsec$ aperture; see Paper I), the overdensity in source galaxy number counts $\delta_{g}$, and the same after randomizing the galaxy positions within the survey mask. All maps are shown at $\texttt{NSIDE}=1024$. While the DECADE magnitude limit map has significantly more structure relative to that of DES Y3, the DECADE source-galaxy sample has much less inhomogeneity as the DECADE and DES Y3 source-galaxy samples are defined with certain cuts (specifically $m_{i}<23.5$, where $m_{i}$ is the $i$-band magnitude of the source galaxy) that make the variation in $\delta_{g}$ similar across the two surveys. The remaining large-scale variations, found towards the eastern/western edges of the footprint, are due to the closer proximity of the DECADE footprint to the Galactic plane. The right-most panels show, as a simple reference, the number-density fluctuations for a sample with randomized positions. There are still some small fluctuations due to the mask structure on small scales; the mask is defined at $\texttt{NSIDE}=4096$ (Paper I).
@@
Our modeling pipeline follows the approach of Krause et al. (2021), who detailed the pipeline for the DES Y3 analyses. We reproduce the salient parts of the model here, and direct the interested reader to that work for a more detailed discussion. We use the two-point correlation function of galaxy shapes, $\xi_{\pm}(\theta)$ (Bartelmann & Schneider 2001), computed for different angular separations $\theta$, as our primary cosmological measurement. We have
@@
$\times\left[C_{EE}^{ij}(\ell)\pm C_{BB}^{ij}(\ell)\right],$ (1)
where $\ell$ is the harmonic multipole, the functions $G^{\pm}_{\ell}(x)$ are computed from Legendre polynomials $P_{\ell}(x)$ and averaged over angular bins (Krause et al. 2021), and the $i$ and $j$ indices specify the two tomographic redshift bins from which the correlation function is calculated. Assuming the extended Limber approximation (Limber 1953; Loverde & Afshordi 2008) and in a spatially flat universe, the lensing power spectrum can be written as
$C^{ij}_{EE}(\ell)=\int_{0}^{\chi_{\rm H}}d\chi\frac{q^{i}(\chi)q^{j}(\chi)}{\chi^{2}}P_{\rm NL}\left(\frac{\ell+1/2}{\chi},\chi\right),$ (2)
where $\chi$ is the radial comoving distance, $\chi_{\rm H}$ is the distance to the horizon, $P_{\text{NL}}$ is the nonlinear matter power spectrum, and $q(\chi)$ is the lensing efficiency defined via
$q^{i}(\chi)=\frac{3}{2}\Omega_{\rm m}\left(\frac{H_{0}}{c}\right)^{2}\frac{\chi}{a(\chi)}\int_{\chi}^{\chi_{H}}d\chi^{\prime}n_{i}(\chi^{\prime})\frac{\chi^{\prime}-\chi}{\chi^{\prime}},$ (3)
where $\Omega_{\rm m}$ is the matter density today, $H_{0}$ is the Hubble parameter today, $a$ is the scale factor, $c$ is the speed of light, and $n_{i}(\chi)$ is the redshift distribution of the bin $i$. The $B$-mode angular power spectrum, $C^{ij}_{BB}$, will be non-zero due to the presence of intrinsic alignments (IA) of galaxy shapes (e.g., Troxel & Ishak 2015; Blazek et al. 2019); [^2] In practice, lensing also generates a $C^{ij}_{BB}$ term but its amplitude is orders of magnitude lower than the $C^{ij}_{EE}$ term and is therefore ignored (e.g., Krause & Hirata 2010). see Section 3.2 for discussions on our IA model.
@@ 3.1 Nonlinear power spectrum
On small scales, the nonlinearity of structure formation as induced by gravity necessitates a modification of the matter power spectrum from the linear model, where the latter is typically calculated via a Boltzmann solver such as CAMB (Lewis et al. 2000) or Class (Lesgourgues 2011). As the resolution and fidelity of $N$-body simulations have improved, so too have prescriptions for the nonlinear power spectrum. In the DES Y1 and Y3 cosmic shear analysis, HaloFit (Takahashi et al. 2012) was used and validated as the default nonlinear power spectrum. However, the study and joint-analysis between DES and the Kilo Degree Survey (KiDS) in DES and KiDS Collaborations et al. (2023) show that the updated version of HMCode (HMCode2020, Mead et al. 2020, 2021) is more consistent with the $N$-body simulation-based Euclid emulator (Euclid Collaboration: Knabenhans et al. 2019). At $k<10h{\rm Mpc}^{-1}$, HMCode2020 has an accuracy of $2.5\%$ when compared to $N$-body simulations. As a result, we adopt HMCode2020 as the fiducial model for our nonlinear power spectrum. Note that for the constraining power of the DECADE data, we do not expect any significant differences in constraints when switching between the HaloFit or HMCode2020 models. We explicitly test this modeling choice in our simulated chains below, and also check it directly on the data in Paper IV.
@@
The above discussion pertains to just the impact of gravitationally sourced nonlinear evolution in the matter power spectrum. The structure on small scales is also impacted by the non-gravitational processes of baryons; most prominently, the ejection of gas from the halo due to feedback processes (Chisari et al. 2018). Hydrodynamical simulations (see Vogelsberger et al. 2020, for a review), which model these effects through approximate subgrid prescriptions, generate a variety of predictions for the properties of halos in simulations (e.g., Anbajagane et al. 2020; Lim et al. 2021; Lee et al. 2022; Stiskalek et al. 2022; Anbajagane et al. 2022a, b; Shao et al. 2023; Shao & Anbajagane 2024; Gebhardt et al. 2023) and therefore also for the changes in the nonlinear matter power spectrum (e.g., van Daalen et al. 2011; Springel et al. 2018; Chisari et al. 2018; Amon & Efstathiou 2022; Salcido et al. 2023). These variations have been captured through phenomenological, halo-based models such as “baryonification” (Schneider & Teyssier 2015; Schneider et al. 2019; Aricò et al. 2021; Anbajagane et al. 2024a) and also a variant of the HMCode2020 model discussed above, denoted as “HMCode2020 Feedback” (following the nomenclature used in CAMB).
@@
The “HMCode2020 Feedback” model includes a $T_{\rm AGN}$ parameter (here, AGN stands for Active Galactic Nuclei) — which can be interpreted as an “effective feedback” parameter with larger values indicating stronger suppression of matter clustering — that quantifies the baryonic contribution to the nonlinear matter power spectrum, and this parameter was varied in the joint-analysis of DES and KiDS (DES and KiDS Collaborations et al. 2023). In this work, our fiducial analysis does not use any of these techniques and follows DES Y3 (Secco & Samuroff et al. 2022; Amon et al. 2022): we remove any scales where the baryonic effects are significant and analyze the remaining scales with a model that does not account for baryons. The identification of scales to be removed is done using the procedure described in Section 4.1. However, we also perform a variant analysis in Paper IV that uses all measurements (i.e., does not perform any scale cuts) and instead varies the $T_{\rm AGN}$ parameter to include the effects of baryons.
@@ 3.2 Intrinsic alignment (IA)
While cosmic shear manifests in the spatial correlations of galaxy shapes, these shapes can also be spatially correlated in the absence of lensing due to the gravitational environment, or tidal fields, that the galaxies form within (see, e.g., the reviews of Troxel & Ishak 2015; Kiessling et al. 2015; Lamman et al. 2024, and references therein). This effect is referred to as intrinsic alignment. The IA signal is directly connected to the matter distribution and therefore can be predicted using the matter power spectrum.
The DES Y3 cosmic shear analysis uses the Tidal Alignment and Tidal Torquing model (TATT, Blazek et al. 2019), which allows for five free nuisance parameters: a tidal alignment amplitude ($a_{1}$) with redshift evolution ($\eta_{1}$); a tidal torquing amplitude ($a_{2}$) with redshift evolution ($\eta_{2}$); and a linear bias amplitude ($b_{\rm TA}$). The TATT model reduces to the simpler, Nonlinear linear alignment model (NLA, Hirata & Seljak 2004; Bridle & King 2007; Joachimi et al. 2011) when setting $a_{1}=A_{\rm IA}$, $\eta_{1}=\eta$, and $a_{2}=\eta_{2}=b_{\rm TA}=0$. Different analyses of DES Y3 have used either TATT (Secco & Samuroff et al. 2022; Amon et al. 2022; Samuroff et al. 2023) or NLA (DES and KiDS Collaborations et al. 2023) as their fiducial choice. For the DECADE cosmic shear analysis, we adopt the more conservative TATT model as our fiducial choice but test the change in our results if we revert to the simpler NLA model. [^3] Since TATT has more parameters, it may allow the fit to explore unphysical or weakly constrained regions of parameter space. Thus, results are often cross-checked using the simpler NLA model.
The amplitude of the IA contribution is parameterized as,
$A_{1}(z)$ $=-a_{1}\bar{C}_{1}\frac{\rho_{\text{crit}}\Omega_{\rm m}}{D(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{1}},$ (4)
$A_{2}(z)$ $=5a_{2}\bar{C}_{1}\frac{\rho_{\text{crit}}\Omega_{\rm m}}{D^{2}(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{2}},$ (5)
$A_{1\delta}(z)$ $=b_{\rm TA}A_{1}(z),$ (6)
where $A_{1}$ and $A_{2}$ scale the matter power spectra, $D(z)$ is the linear growth rate, $\rho_{\text{crit}}$ is the critical density at $z=0$, and $\bar{C}_{1}=5\times 10^{-14}M_{\odot}h^{-2}\text{Mpc}^{2}$ is a normalization constant, set by convention. We choose a pivot redshift, $z_{0}=0.62$ following Secco & Samuroff et al. (2022); Amon et al. (2022). The free parameters of our model are the amplitudes $a_{1},a_{2},b_{\rm TA}$ and the power-law indices $\eta_{1},\eta_{2}$. As mentioned above, the NLA model is obtained by setting $a_{2},\eta_{2},b_{\rm TA}=0$. See Equations 20–23 in Secco & Samuroff et al. (2022) for a description of the different IA-related power-spectra that contribute to the final signal.